simplify
2 * the square root of 3/the squar root of 48

Answer by RAY100(1637)

( 2*sqrt3) / sqrt48 = (2 * sqrt 3) / (sqrt 3*sqrt 16 ) =2/sqrt16 =2/4 =1/2